hi, i really dont understand some things in algebra. on my worksheet it says to graph this linear equation using three points. the equation is 2y-2=6x
please help me, and please explain how you did this, and if possible please tell me the answer before monday so i can do good on this worksheet.
thanks in advance.

Answer by music974(43)

The first thing you should do on this equation is change it so that y is equal to something. 2y-2=6x first you add 2 from each side because you must do the inverse operation of -2, this gives you 2y=6x+2.Now you can solve the equation. You can use x=0, x=1, x=2. 2y=6(0)+2 / 2y=6(1)+2 / 2y=6(2)+2
2y=0+2 / 2y=6+2 / 2y=12+2
2y=2 / 2y=8 / 2y=14
y=1 / y=4 / y=7

The three points would be (0,1),(1,4),(2,7).